### 18 Mar 2026 v2

# Power-Law Running of the Higgs Mass

Abstract
The renormalized scalar mass squared is a function of the energy scale and power-runs as its square: it is $\mathcal{O}(M_{\text{GUT}}^{2})$ at $M_{\text{GUT}}^{2}$ and $\mathcal{O}(m_{W}^{2})$ at $m_{W}^{2}$. There is no hierarchy problem.
The Higgs boson mass parameter ATLAS:Higgs:2012 ; CMS:Higgs:2012 $m_{h}^{2}$, close to $(125~\text{GeV})^{2}$, is 28 orders of magnitude below the square of the scale $M_{\text{GUT}}\sim 10^{16}$ GeV at which the Standard Model is expected to be completed by new physics such as string theory, not necessarily Grand Unified Theory (GUT). In the Standard Model (SM), the one-loop correction to the scalar mass squared from a field of mass $M_{\text{GUT}}$ takes the form $\delta m_{h}^{2}\sim g^{2}M_{\text{GUT}}^{2}$, where $g$ is the relevant coupling. The correction overwhelms the observed mass, requiring an apparent cancellation between huge quantities—the bare mass and the loop correction—to one part in $10^{28}$.
This is the hierarchy problem Veltman:1981 ; tHooft:1980 ; Susskind:1979 ; Barbieri:Giudice:1988 ; Giudice:2008 . ’t Hooft gave the symmetry account tHooft:1980 : a mass parameter is protected if setting it to zero increases the symmetry. For fermions, setting the mass to zero, $m_{f}\to 0$, restores chiral symmetry; for the scalar, $m_{h}\to 0$ restores no symmetry. The resulting corrections are
$\delta m_{f}$ $\propto m_{f}\log M_{\text{GUT}}^{2}$ fermion (protected), (1)
$\delta m_{h}^{2}$ $\propto M_{\text{GUT}}^{2}$ scalar (unprotected).
This was taken to mean that a light scalar is unnatural unless a new symmetry (supersymmetry Dimopoulos:Georgi:1981 , compositeness Kaplan:Georgi:1984 , extra dimensions ArkaniHamed:1998 ; Randall:Sundrum:1999 ) stabilizes it, or the value is environmentally selected Bousso:Polchinski:2000 .
The Large Hadron Collider has found no evidence for such symmetries. We show that none are needed.
The fully renormalized mass function virtual ; Choi:2023:observables ; Choi:2025:Higgs shows that the scalar mass squared runs as the square of the energy scale $q^{2}$. This is not an additional assumption but a consequence of the standard quantum field theory tHooft:1971 ; tHooft:Veltman:1972 , once the renormalization is carried out completely. Therefore it is $\mathcal{O}(M_{\text{GUT}}^{2})$ at $M_{\text{GUT}}^{2}$ and $\mathcal{O}(m_{W}^{2})$ at $m_{W}^{2}$—the natural scaling of an unprotected quantity. In a grand unified theory, an order-one boundary condition at the unification scale suffices to produce an electroweak-scale Higgs, replacing the traditional $10^{-28}$ cancellation.
The $q^{2}$-dependence of the scalar self-energy has long been perceived—but only as a part of the divergent, unrenormalized expression. It appeared in the earliest one-loop calculations Veltman:1981 ; Fleischer:Jegerlehner:1981 ; gauge-invariant formulations were developed in Cornwall:1982 ; Binosi:Papavassiliou:2009 ; the physical significance of the quadratic term was explored in Jegerlehner:2014 ; Jegerlehner:2021 ; and the Wilsonian functional renormalization group Wilson:Kogut:1974 , which naturally captures the quadratic scaling, was recently applied to the full SM in Garces:2025 . These works treat the $q^{2}$ dependence within the framework of cutoff-dependent bare parameters or Wilsonian flows. What is new in Choi:2023:observables ; Choi:2025:Higgs is the extraction of a finite, physical mass function in which the $q^{2}$ dependence is not a cutoff artifact or a source of fine-tuning but the entire observable content of the scalar mass.
The mass function
The mass function constructed in Ref. virtual defines the scalar mass at every external momentum $q$. On-shell renormalization yields Choi:2023:observables ; Choi:2025:Higgs
$\begin{split}m^{2}(q^{2})&=m_{h}^{2}+\Sigma(q^{2})-\Sigma(m_{h}^{2})-(q^{2}-m_{h}^{2})\frac{d\Sigma}{dq^{2}}(m_{h}^{2})\\ &\equiv m_{h}^{2}+\Sigma_{\text{ren}}(q^{2}),\end{split}$ (2)
with the dressed propagator $i/(q^{2}-m^{2}(q^{2}))$ and pole mass $m_{h}$ fixed by $m^{2}(m_{h}^{2})=m_{h}^{2}$. [^1] Above particle-production thresholds, $\Sigma$ acquires an imaginary part; $\Sigma_{\text{ren}}$ denotes $\text{Re}\,\Sigma_{\text{ren}}$ throughout
The mass function is finite Bogoliubov:Parasiuk:1957 ; Hepp:1966 ; Zimmermann:1969 , and regularization-independent Choi:2023:observables ; Choi:2024:decoupling . In particular, the $q^{2}$ dependence of $m^{2}(q^{2})$ is a physical effect that is present regardless of the renormalization scheme. The contribution from gauge bosons can be made gauge-parameter invariant by absorbing non-invariant terms to the vertices, thanks to the Ward–Takahashi identity. The result coincides with the Feynman gauge choice, which we assume from now on Cornwall:1982 ; Binosi:Papavassiliou:2009 ; Papavassiliou:1997fn ; Papavassiliou:1997pb ; virtual .
What runs is the mass function $m^{2}(q^{2})$, the full momentum-dependent dressing. The pole mass is a fixed number; it does not run. The power-law behavior at momenta far from the pole is a prediction; the pole mass itself is data.
The mass function $m^{2}(q^{2})$ is the same kind of object as the running coupling $\alpha_{s}(q^{2})$ Gross:Wilczek:1973 ; Politzer:1973 : both are gauge-invariant, physically measurable, and defined at the recoil momentum of the process.
Two properties of the mass function are central to what follows.
  1. 1.
    Decoupling The complete renormalization (2) removes the leading sensitivity to heavy fields. A field of mass $M^{2}\gg q^{2}$ contributes Choi:2024:decoupling ; Choi:2024:stability
    $\Sigma_{\text{ren}}^{\text{heavy}}(q^{2})={\cal O}\left(\frac{(q^{2}-m_{h}^{2})^{2}}{M^{2}}\right),$ (3)
    extending the Appelquist–Carazzone theorem Appelquist:Carazzone:1975 to the super-renormalizable scalar mass. In (3), $m_{h}$ and $M$ are both pole masses—physical, measurable quantities—so the suppression is a physical statement, independent of any scheme or convention.
  2. 2.
[…]
    $\Sigma_{\text{ren}}(q^{2})$ $\sim q^{2}\log q^{2}$ scalar,
    This verifies the ’t Hooft account (1). The crucial difference is that the ’t Hooft estimate gives a correction at $M_{\text{GUT}}$, whereas the renormalized mass function $m^{2}(q^{2})$ is defined at every scale and is computed bottom-up from the measured pole mass. Using $m_{h}=125.08$ GeV ATLAS:2023 ; CMS:2025 , the complete one-loop correction to the Higgs mass parameter (see Apeendix) gives, at $\sqrt{q^{2}}=M_{\text{GUT}}=10^{16}$ GeV,
    $\frac{m^{2}(M_{\text{GUT}}^{2})}{M_{\text{GUT}}^{2}}\approx\underbrace{1.135}_{\text{top}}\underbrace{{}-0.169}_{W}\underbrace{{}-0.109}_{Z}+\underbrace{0.002}_{h}=+0.858,$ (5)
    which is $\mathcal{O}(1)$. The mass function, measured in units of the scale at which it is probed, is order one at every scale.
Decoupling resolves the technical part of the hierarchy problem, which results in inconsistency Choi:2024:decoupling ; Choi:2024:stability . Super-renormalizable quantities like the Higgs mass have been suspected of violating decoupling, giving $\delta m_{h}^{2}\sim M^{2}$. This expectation follows from naive dimensional analysis of the unrenormalized self-energy; the complete renormalization (2) replaces it with (3).
The 28 orders of magnitude between $m_{h}^{2}$ and $M_{\text{GUT}}^{2}$ are the natural scaling of a dimension-2 quantity that runs as $q^{2}$—not a hierarchy that requires explanation. The reason is nothing but the absence of symmetry protection of the scalar mass.
Power-law running of the scalar
The total can be written as ($v=246$ GeV is the Higgs vacuum expectation value)
$\frac{\Sigma_{\text{ren}}^{\text{SM}}}{q^{2}}=\frac{N_{c}m_{t}^{2}}{8\pi^{2}v^{2}}\log\frac{q^{2}}{m_{t}^{2}}-\cdots\approx 0.0142\log\frac{q^{2}}{m_{\text{EW}}^{2}}-0.045,$ (6)
where $m_{\text{EW}}$ denotes a representative electroweak mass ($m_{t}$, $m_{W}$, $m_{Z}$ are all $\mathcal{O}(m_{\text{EW}})$; threshold differences are absorbed into the constant). The leading top coefficient $N_{c}m_{t}^{2}/(8\pi^{2}v^{2})=0.0187$ ($N_{c}=3$ is the number of quark colors) is partially cancelled by the $W$ (0.0027-0.0027) and $Z$ (0.0017-0.0017) contributions, giving the net slope $0.0142$. The last term is the sum of the constant terms from the large-$q^{2}$ limits (e.g., 4/3-4/3 from the top in Eq. (14), plus the $W$ and $Z$ constants). The Higgs self-coupling also runs as a power law (see Eq. (13)) with coefficient $\sim m_{h}^{2}/(32\pi^{2}v^{2})$, but it is numerically negligible (+0.002+0.002 at $M_{\text{GUT}}$). The coefficient is $\mathcal{O}(10^{-2})$ but multiplies $\log(M_{\text{GUT}}^{2}/m_{\text{EW}}^{2})\approx 63$, giving an order-one product.
In the $\overline{\text{MS}}$ scheme Bardeen:Buras:Duke:Muta:1978 with dimensional regularization, quadratic powers of the renormalization scale $\mu^{2}$ are never generated—they correspond to poles at $d=2$, which are invisible at $d=4-\epsilon$. The beta function for $\bar{m}^{2}(\mu)$ is therefore purely logarithmic, and the scalar mass appears to run like a fermion mass. With no $\mu^{2}$ term in sight, there is nothing to trigger alarm—and the power-law $q^{2}$ dependence, which lives in the finite parts of the amplitude, goes unrecognized. This does not mean that dimensional regularization is at fault—the mass function $m^{2}(q^{2})$ is regularization-independent, as noted above—but that $\overline{\text{MS}}$ is a scheme in which the physical scale dependence remains implicit.
The validity of these results requires that the perturbation theory remain under control at all scales up to $M_{\text{GUT}}$ Degrassi:2012 ; Buttazzo:2013 . The one-loop correction $\Sigma_{\text{ren}}^{\text{SM}}/q^{2}\approx 0.86$ at $M_{\text{GUT}}$ is less than unity, so the propagator denominator $q^{2}-m^{2}(q^{2})=q^{2}(1-0.86)>0$ remains positive—no sign change, no ghost, no breakdown. The top quark alone gives +1.135+1.135, which exceeds unity, but the $W$ and $Z$ contributions (0.28-0.28 combined) bring the total below one; the full SM content is essential (see Appendix). This partial cancellation is not accidental: fermion loops carry a factor (1)(-1) from the closed Grassmann trace (spin-statistics theorem Pauli:1940 ), giving the top a positive contribution while the $W$ and $Z$ contribute with opposite sign. We have verified that the sign structure persists at two loops—each fermion loop carries its (1)(-1)—so the partial cancellation is a structural property of the SM’s spin content, not a one-loop coincidence. The two-loop correction is suppressed by a further factor of $m_{\text{EW}}^{2}/(16\pi^{2}v^{2})\log(M_{\text{GUT}}^{2}/m_{\text{EW}}^{2})$, giving $\Sigma^{(2)}/\Sigma^{(1)}\sim 0.1$$0.2$; this shifts the total from $0.86$ to roughly 0.90.91.01.0, which does not change the order-one character of $m^{2}(M_{\text{GUT}}^{2})/M_{\text{GUT}}^{2}$. That $r$ approaches unity at $M_{\text{GUT}}$ is expected: this is the boundary of the SM’s perturbative domain, where the GUT description takes over. Perturbation theory is controlled at every scale from the electroweak to the GUT.
Constraint on GUT
We now examine what this implies for a UV completion. The Standard Model must be completed at some high scale, whether by grand unification, string theory, or another framework. The scale $M_{\text{GUT}}\sim 10^{16}$ GeV indicated by gauge coupling unification Georgi:Quinn:Weinberg:1974 ; Langacker:Luo:1991 is representative of where such new physics is expected to appear; in $SU(5)$, the $X$ boson pole mass provides a more precise candidate for $M_{\text{GUT}}$. We take $SU(5)$ grand unification Georgi:Glashow:1974 as a concrete prototype; the specific parameters differ between UV completions, but the essential structure—an order-one boundary condition at the high scale, with the SM running below it—does not.
Below $M_{\text{GUT}}$, only SM fields run. The result (5) is computed within the Standard Model, using $m_{h}=125$ GeV as input. It gives the SM side of the matching: $r\approx 0.86$, fully computed from measured parameters. No GUT parameters enter.
The ratio $\Sigma_{\text{ren}}^{\text{SM}}/q^{2}$ grows logarithmically $\sim\log q^{2}$ and reaches unity near $10^{18}$ GeV, providing an upper bound on the SM’s domain of validity—in analogy with the Landau pole in the $U(1)_{Y}$ coupling, which independently requires new physics below that scale.
Above $M_{\text{GUT}}$, the GUT dynamics determines the Higgs doublet mass at $M_{\text{GUT}}$. In $SU(5)$, the boundary condition is
$r\equiv\frac{\mu^{2}-\frac{3}{2}\lambda_{5}V+\Sigma^{\text{heavy}}_{\text{GUT}}(M_{\text{GUT}}^{2})}{M_{\text{GUT}}^{2}},$ (7)
where $\mu^{2}$ is the $\mathbf{5}$-Higgs mass parameter, $\lambda_{5}$ the coupling to the adjoint, $V$ the adjoint VEV, and $\Sigma^{\text{heavy}}_{\text{GUT}}$ contains loops from $X$, $Y$ bosons, the triplet, and the adjoint. This is the matching and the value $r$ provides a constraint on the GUT.
The traditional formulation begins from a speculative UV scale, postulates a bare Lagrangian there, and extrapolates downward using an unrenormalized self-energy that violates decoupling—arriving at a “crisis” constructed entirely from unobservable quantities. The problem was not in the physics but in the direction of the argument.
Explaining the electroweak scale
Figure 1: The SM boundary condition $r=m^{2}(M_{\text{GUT}}^{2})/M_{\text{GUT}}^{2}$ vs. the pole mass $m_{h}$. The range $r\approx 0.57$$0.89$ maps onto $m_{h}=50$$2000$ GeV. The point corresponds to the observed value. Cusps at $m_{h}=2m_{W}$, $2m_{Z}$, and $2m_{t}$ are physical threshold effects.
The matching also runs in the other direction. Given GUT parameters, matching equates the GUT boundary condition (7) to the SM mass function:
$r=\frac{m_{h}^{2}+\Sigma^{\text{SM}}_{\text{ren}}(M_{\text{GUT}}^{2})}{M_{\text{GUT}}^{2}},$ (8)
and the SM running then predicts $m_{h}$. Inverting Eq. (8) for $m_{h}$ gives
$m_{h}^{2}=rM_{\text{GUT}}^{2}-\Sigma_{\text{ren}}^{\text{SM}}(M_{\text{GUT}}^{2}).$ (9)
The difference is tiny but no longer fine-tuned. Given $M_{\text{GUT}}^{2}$ and $r$ from the GUT, $\Sigma_{\text{ren}}^{\text{SM}}$ is fully determined by the SM running. The top-down construction determines the entire running from $M_{\text{GUT}}^{2}$ to $m_{h}^{2}$.
For benchmarking, we vary the Higgs pole mass over $m_{h}=50$$2000$ GeV. Fig. 1 shows the SM boundary condition $r$ as a function of $m_{h}$. The ratio is $\mathcal{O}(1)$ for any electroweak-scale pole mass, varying by only $\sim 0.3$ on that range. This is natural dimensional scaling: $m^{2}$ has mass dimension 2 and runs as $q^{2}$, so $m^{2}(q^{2})/q^{2}$ is $\mathcal{O}(1)$ at every scale.
The constraint on the GUT (7) is
$r\in[0.57,0.86],$ (10)
an order-one requirement on $\mathcal{O}(1)$ parameters. In the traditional framework, the same constraint was an accuracy of $\mathcal{O}(10^{-28})$. The target has widened to a band of width $\sim 0.3$. The specific value $m_{h}=125$ GeV within the electroweak band is a precise constraint on the GUT—analogous to the constraint that the top Yukawa is $0.99$ rather than 0.50.5—but it is a constraint on $\mathcal{O}(1)$ parameters, not a fine-tuning.
Conversely, if the GUT delivers an order-one $r$, inverting the relation gives the prediction $m_{h}(r)$. For a sizable range of order-one $r$, the result is naturally at the electroweak scale.
The same mechanism suggests a possible resolution of the doublet-triplet splitting problem Sakai:1981 ; Dimopoulos:Wilczek:1981 . Both the doublet and triplet mass functions are $\mathcal{O}(M_{\text{GUT}}^{2})$ at the GUT scale. Below $M_{\text{GUT}}$, the triplet mass function also runs as $q^{2}$, but with different coefficients determined by its color-triplet quantum numbers. Because these are fixed by the gauge group, the running is model-independent: it applies to any GUT containing a color-triplet scalar. If the triplet’s running is sufficiently slow that its pole mass remains at $\mathcal{O}(M_{\text{GUT}})$, no special splitting mechanism would be needed. The explicit calculation is left to future work.
From naturalness to flavor
The remaining question—why does the GUT deliver $r\in[0.57,0.86]$?—is not a naturalness problem requiring new symmetries at the TeV scale. It is a question about the specific value of an order-one coupling, not protected by any symmetry, yet never regarded as unnatural. Even the electron Yukawa $y_{e}\sim 10^{-6}$, six orders of magnitude below the top Yukawa $y_{t}\approx 1$ and far more “unexplained” than any $\mathcal{O}(1)$ ratio, is not considered a naturalness crisis. The specificity of the GUT-scale ratio belongs to the same category of questions about coupling values, approachable through whatever “flavor” symmetry organizes the couplings.
To sum up, the power-law running makes the scalar mass $m^{2}(q^{2})$ scale as $q^{2}$, so $m^{2}(q^{2})/q^{2}$ is $\mathcal{O}(1)$ at every scale. Any order-one GUT-scale boundary condition yields an electroweak-scale pole mass through the SM running. If the scalar mass ran logarithmically like a fermion mass, $m^{2}$ would barely change between scales, and $(125~\text{GeV})^{2}$ from $\mathcal{O}(M_{\text{GUT}}^{2})$ inputs would require $10^{-28}$—the very fine-tuning that the hierarchy problem describes.
The absence of protection is not the disease. It provides an explanation of the hierarchy. An experimental test through off-shell Higgs production Kauer:Passarino:2012 ; Caola:Melnikov:2013 at the LHC is discussed in Ref. Choi:2025:Higgs .
In fact, the same argument applies to the cosmological constant $\Lambda$ with the mass dimension four. In flat space, it is a constant. In curved space, it couples to the graviton via $\sqrt{\det g}\Lambda$ coupling, with which $\Lambda(q^{2})$ acquires the external momentum dependence $q^{2}$. The same power-law applies to $\Lambda(q^{2})/q^{4}={\cal O}(1)$ at every scale $q^{2}.$ We call for the urgent renormalization of quantum gravity.
Appendix A Appendix: Explicit one-loop formulas
We collect the renormalized one-loop self-energies from Ref. Choi:2025:Higgs that yield the result (5). The total mass function is $m^{2}(q^{2})=m_{h}^{2}+\Sigma_{\text{ren}}^{\text{SM}}(q^{2})$, where $\Sigma_{\text{ren}}^{\text{SM}}=\sum_{a=t,W,Z,h}\text{Re}\Sigma_{1,\text{ren}}^{a}(q^{2})$, with $\Delta_{m}(q^{2})\equiv m^{2}-x(1{-}x)q^{2}$, where $x$ is the Feynman parameter and $m$ is the mass of the particle in the loop, and $L_{M}(q^{2})\equiv m_{h}^{4}-2q^{2}(1{+}3x^{2})M^{2}+10M^{4}$.
Top quark:
$\Sigma_{1,\text{ren}}^{t}(q^{2})=\frac{3m_{t}^{2}N_{c}}{4\pi^{2}v^{2}}\int_{0}^{1}\!\!dx\biggl[\Delta_{m_{t}}(q^{2})\log\frac{\Delta_{m_{t}}(m_{h}^{2})}{\Delta_{m_{t}}(q^{2})}\\ -(q^{2}-m_{h}^{2})x(1{-}x)\biggr].$ (11)
$W$ boson (transverse + Goldstone + ghost):
$\Sigma_{1,\text{ren}}^{W}(q^{2})=\frac{-1}{16\pi^{2}v^{2}}\int_{0}^{1}\!\!dx\biggl[L_{m_{W}}(q^{2})\log\frac{\Delta_{m_{W}}(m_{h}^{2})}{\Delta_{m_{W}}(q^{2})}\\ -(q^{2}-m_{h}^{2})\frac{x(1{-}x)L_{m_{W}}(m_{h}^{2})}{\Delta_{m_{W}}(m_{h}^{2})}\biggr].$ (12)
$Z$ boson: Same structure with an overall factor $1/2$ and $m_{W}\to m_{Z}$.
Higgs self-coupling:
$\Sigma_{1,\text{ren}}^{h}(q^{2})=\frac{-9m_{h}^{4}}{32\pi^{2}v^{2}}\int_{0}^{1}\!\!dx\biggl[\log\frac{\Delta_{m_{h}}(m_{h}^{2})}{\Delta_{m_{h}}(q^{2})}\\ -(q^{2}-m_{h}^{2})\frac{x(1{-}x)}{\Delta_{m_{h}}(m_{h}^{2})}\biggr].$ (13)
At $q^{2}\gg m_{t}^{2}$, the top contribution simplifies to
$\frac{\Sigma_{1,\text{ren}}^{t}}{q^{2}}=\frac{N_{c}m_{t}^{2}}{4\pi^{2}v^{2}}\biggl[\frac{1}{2}\log\frac{q^{2}}{m_{t}^{2}}-\frac{4}{3}\biggr],$ (14)
exhibiting the characteristic $q^{2}\log q^{2}$ power-law behavior: the self-energy grows as $q^{2}$ times a logarithm, so $\Sigma/q^{2}$ grows logarithmically—but the mass function $m^{2}(q^{2})\approx m_{h}^{2}+\Sigma_{\text{ren}}(q^{2})$ grows as $q^{2}$ itself. The $W$ and $Z$ contributions have the same power-law structure with opposite sign:
$\frac{\Sigma_{1,\text{ren}}^{W}}{q^{2}}\approx-\frac{m_{W}^{2}}{4\pi^{2}v^{2}}\log\frac{q^{2}}{m_{W}^{2}}+\text{const},$ (15)
and similarly for $Z$ with a factor $1/2$. The numerical evaluation of these integrals at $\sqrt{q^{2}}=10^{16}$ GeV, using $m_{t}=172.56$ GeV, $m_{W}=80.37$ GeV, $m_{Z}=91.19$ GeV, $v=246.26$ GeV PDG:2024 , yields the sector-by-sector values quoted in Eq. (5).
Acknowledgements.
The author is grateful to Jaehoon Jeong, Taehyun Jung, Hyun-Min Lee and Matthew McCollough for discussions. This work is partly supported by the grant RS-2023-00277184 of the National Research Foundation of Korea. The author used Claude (Anthropic) for editing and discussion during the manuscript preparation. The final content was reviewed and approved by the author, who takes full responsibility for the work.